A discrete uniform population consists of the values 0,1,2, and 3. Suppose you have a sample size of 2 without replacement, find population and sample mean, variance, and standard deviation.


1
Expert's answer
2021-05-24T16:08:15-0400

"Number = \frac{N!}{n!(N-n)!} \\

= \frac{4!}{2!(4-2)!} \\

= \frac{3 \times 4}{2} \\

= 6"

samples:

0, 1

0, 2

0, 3

1, 2

1, 3

2, 3

Population mean "= \frac{0+1+2+3}{4}=1.5"

Population variance "= \frac{1}{4-1}( (0-1.5)^2+(1-1.5)^2 +(2-1.5)^2 +(3-1.5)^2 )"

"= \frac{1}{3}(2.25+0.25+0.25+2.25) = 1.667"

Population standard deviation "= \sqrt{1.667}=1.29"